• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 22:40:57 on March 02, 2009

    Entry number 264246

    keyword=BigBite DAQ fixed

    Shift crew noted that the DAQ didn't appear to be taking any actual data triggers. No errors, no other problems, scaler events were arriving (so the 'Events this run' box in CODA was incrementing (albeit slowly).

    Turned out that some garbage had gotten typed into the top of the prescale.dat file late last night (after beam had gone away). The TS couldn't parse it and set all prescales to zero (ie. disabled all triggers). The end-of-run scripts DID parse it correctly so normal prescale values did show up in the halog entries.

    Fixed now.

    A copy of this log entry has been emailed to: kalyan